OpenStudy (anonymous):

What is the inverse of y=3^(-x) ?

OpenStudy (bibby):

replace the x and y, solve for y \[\large x=3^{(-y)}\] now what?

OpenStudy (anonymous):

i'm not sure how to go about this after that step. I do know the properties though a^y= x and log basea x = y just not how to apply it here.

OpenStudy (bibby):

exactly. I think if you took the ln of both sides, you'd get something like lnx = -yln3 Then you can divide and multiply by -1

OpenStudy (anonymous):

The answer sheet says it should be y= - logbase3 x, however after doing it the way you suggested I get y= - (lnx)/ (ln3). Does that mean - (lnx)/ (ln3)= - logbase3 x or am I missing a step?

OpenStudy (bibby):

I used ln and not logs, let me try again with one of these

OpenStudy (bibby):

x=3^(−y) -y*log3=logx y= -log x ----- log 3 is that what you get?

OpenStudy (bibby):

\[x=3^{(−y)}\] \[ -y*\log3=logx\] \[y=\frac{ -\log x }{ \log 3 }\]

OpenStudy (anonymous):

Yeah I get that in ln as well, but I was wondering if it's the same as y= - logbase3 x since it's in the answer sheet.

OpenStudy (jdoe0001):

\(\bf y=log_3x\Longleftarrow \textit{using change of base rule }\implies y=\cfrac{log(x)}{log(3)}\)

OpenStudy (jdoe0001):

\(\bf y=-log_3x\Longleftarrow \textit{using change of base rule }\implies y=\cfrac{-log(x)}{log(3)}\) anyhow, yes, it's the same :)

OpenStudy (bibby):

I don't even remember that rule lol, thanks jdoe

OpenStudy (jdoe0001):

@bibby \(\bf \textit{using change of base rule }\implies log_{\color{red}{ a}}{\color{blue}{ b}}\implies \cfrac{log_c{\color{blue}{ b}}}{log_c{\color{red}{ a}}}\) and "c" can be any value, so long it's the same above and below

OpenStudy (anonymous):

Oh ok another question I had is the step when switching 3^(-x) = y for inverse using the property, do I write log base3 y = x or log base3 y= -x ? I'm confused whether the negatives carries or not

OpenStudy (jdoe0001):

ven

OpenStudy (jdoe0001):

acck... anyway Ventricate for the inverse, you do as always, just swap about the variables

OpenStudy (anonymous):

So the negative wouldn't be swapped with the variable right?

OpenStudy (jdoe0001):

right, just the variable itself

OpenStudy (anonymous):

oh ok thank you so much @jdoe0001 and @bibby :)

OpenStudy (jdoe0001):

yw